# DiffHairCard: Auto Hair Card Extraction with Differentiable Rendering

### v1

@@ Abstract.
Hair cards remain a widely used representation for hair modeling in real-time applications, offering a practical trade-off between visual fidelity, memory usage, and performance. However, generating high-quality hair card models remains a challenging and labor-intensive task. This work presents an automated pipeline for converting strand-based hair models into hair card models with a limited number of cards and textures while preserving the hairstyle appearance. Our key idea is a novel differentiable representation where each strand is encoded as a projected 2D curve in the texture space, which enables end-to-end optimization with differentiable rendering while respecting the structures of the hair geometry. Based on this representation, we develop a novel algorithm pipeline, where we first cluster hair strands into initial hair cards and project the strands into the texture space. We then conduct a two-stage optimization, where our first stage optimizes the orientation of each hair card separately, and after strand projection, our second stage conducts joint optimization over the entire hair card model for fine-tuning. Our method is evaluated on a range of hairstyles, including straight, wavy, curly, and coily hairs. To capture the appearance of short or coily hairs, our method comes with the support for hair cap and cross-card.
@@
ccs: Computing methodologies Shape modeling
@@ 1. Introduction
Hair plays a significant role in creating believable characters and immersive environments in films, video games, and virtual reality. However, an average human scalp has about 100,000 - 150,000 hair follicles. Although strand-based hair simulation and rendering have become increasingly popular in recent production pipelines due to their ability to realistically capture the dynamic behavior and visual richness (Tafuri, 2019; Epic Games, 2021; Huang et al., 2023; Hsu et al., 2023), hair cards remain a widely adopted technique in real-time applications (Jiang, 2016). Hair cards are flat, textured quad strips that approximate the visual appearance of hair clusters. As illustrated in Fig. 1, hair cards can effectively reproduce the intricate structure of complex hair styles. Beyond their role in representing high-fidelity hair with a large number of textured planes (Jiang, 2016), hair card representations are also widely used in level-of-detail (LoD), where strand-based models serve as a high-resolution representation, while hair cards can be used as a low-resolution approximation to significantly reduce rendering costs when hair is viewed from a distance.
@@
Unfortunately, the industry currently heavily relies on crafting hairstyles manually using hair cards. In this workflow, artists first create a set of hair card textures that contain varying numbers of strands and degrees of curvature. They then manually extrude each hair card from the scalp outward, following the hair flow while varying its length, width, and orientation to achieve a realistic appearance. Apparently, the fidelity quality of the hair card model is primarily bound by the number of cards and textures allowed. As a result, manually crafting a low-resolution hair card model—with a limited number of quads and card textures while maintaining high visual fidelity to the reference strand-based model—is a challenging and labor-intensive task, typically requiring several days to weeks of work from an experienced artist.
@@
Most recent research works have focused on reconstructing strand-based hair models from single-view images (Zhou et al., 2018; Zhang and Zheng, 2019; Wu et al., 2022; Zheng et al., 2023), multi-view inputs (Kuang et al., 2022; Wu et al., 2024b; Zhou et al., 2024), and CT scans (Shen et al., 2023). However, there is a lack of work on converting these reconstructed strands into hair card representations suitable for real-time applications. Extracting hair cards from strand-based hair models presents several challenges. First, there is a strict budget on both the number of hair cards and on the size of texture, imposed by memory and computational efficiency constraints. Therefore, the extraction process must carefully balance visual fidelity with performance. Second, the hairstyle is highly irregular and chaotic, showing significant variation in strand length, shape, and structure, making it challenging to generate a simplified representation while preserving visual details. Recently, Unreal Engine 5 (Epic Games, 2021) introduced a tool for the automatic generation of hair cards from strand-based models (Epic Games, 2025). However, the quality of its generated results remains insufficient for production use, often failing to preserve the appearance of the target hair strand model when the given hair card target is limited.
@@
This paper presents an automated pipeline that converts strand-based hair models into hair cards. The process begins with hair strand clustering, where strands are grouped based on a hair shape similarity metric. For each cluster, we optimize the orientation of the card geometry to minimize the strand projection error. To improve memory efficiency and runtime performance, we perform second-stage clustering on the hair textures, enabling texture sharing across multiple cards. Finally, we jointly optimize the hair card positions, strand shapes, and hair widths to minimize differences in tangent, depth, and coverage between our output cards and the original strand-based model. To this end, conventional RGB-image texture representations can lead to significant aliasing artifacts. Instead, we propose an explicit hair card representation where each strand is projected into texture space as a 2D curve. These 2D curves serve as an intermediate representation that enables high-quality rendering and supports differentiable optimization. To enhance the visual fidelity for short and coily hairstyles, which is challenging for original hair card representation, our framework also supports baking a hair cap texture and generating a pair of crossed cards per hair cluster, enriching the volumetric appearance of the hair.
@@
We evaluate our pipeline on a diverse range of hairstyles, including straight, wavy, curly, and coily hairs, with a limited number of cards and textures. Experiments show that our automated pipeline surpasses both UE automatic solution and manual-crafted cards.
@@ 2. RELATED WORK
Both hair simulation and rendering are computationally expensive due to the large number of individual strands and their complex interactions. To reduce computational costs, researchers have developed various reduced representations over the years, including 2D strips (commonly referred to as hair cards(Koh and Huang, 2001; Ward et al., 2003), cubic lattice structures(Volino and Magnenat-Thalmann, 2006), short hair strips (Guang and Zhiyong, 2002), and volumetric representations (Wu and Yuksel, 2016; Lee et al., 2019). During rendering, these reduced representations are expanded into full hair using baked textures or procedural functions. Hair cards remain widely used in the gaming industry due to their simplicity and efficiency (Jiang, 2016). For a comprehensive overview of hair rendering and simulation, we refer readers to the course by Bertails et al. (2008). However, creating hair cards for multiple LoDs continues to be a labor-intensive process for artists, representing a significant bottleneck in production pipelines.
@@
Traditionally, artists have used tools like Maya XGen to create strand-based and card-based hair models. However, manually crafting hair remains a labor-intensive process. To simplify hair authoring, Yuksel et al. (2009) introduced hair mesh, a volumetric representation that provides high-level editing tools for artists. Sketch-based input methods (Fu et al., 2007; Shen et al., 2020) have also offered user-friendly interaction capability, allowing artists to design hair directly on the screen. For automatic hair modeling, researchers have explored extracting strand-based geometries directly from images. Earlier methods relied on heuristic-based approaches (Kong and Nakajima, 1998; Paris et al., 2008; Jakob et al., 2009; Hu et al., 2017) or leveraged large 3D hair databases (Hu et al., 2015; Chai et al., 2016; Liang et al., 2018) for guidance. Recently, learning-based techniques have demonstrated accuracy and robustness in reconstructing simple hairstyles from single-view images (Zhou et al., 2018; Zhang and Zheng, 2019; Wu et al., 2022; Zheng et al., 2023) and multi-view inputs (Kuang et al., 2022; Rosu et al., 2022; Sklyarova et al., 2023; Wu et al., 2024b; Takimoto et al., 2024; Zakharov et al., 2024; Zhou et al., 2024). Wu et al. (2024a) present a geometric method to generate highly coiled hair. Beyond explicit modeling and reconstruction, hair synthesis methods enable hairstyle transfer from one to another using feature maps (Wang et al., 2009), further refined with learning-based features (Zhou et al., 2023; Sklyarova et al., 2024; Chen et al., 2024; He et al., 2025). However, these techniques primarily aim to generate explicit strand geometries, which are computationally intensive and unsuitable for real-time applications.
@@
In addition to hair, arbitrarily oriented billboards (or impostors) with textures are widely used for rendering trees and forests (Behrendt et al., 2005) and large-scale scenes (Lall et al., 2018; Hladky et al., 2022). To convert a 3D mesh to billboards, Décoret et al. (2003) proposed a greedy optimization algorithm that approximates the 3D model using a discretized plane parameterization in spherical coordinates. This method was later enhanced by Andujar et al. (2004) to produce better-fitting billboards. Silvennoinen et al. (2014) uses a similar concept to generate sets of planars as occluders. However, as hair presents unique challenges due to the extensive strand number and complex geometry, existing billboard generation methods cannot be directly applied.
@@ 3. Problem Statement and Overview
The input to our method is a strand-based hairstyle model =$\mathcal{H}=\{\mathcal{S}_{i}\}$, where each strand $\mathcal{S}_{i}=\{\mathbf{s}_{i,j}\}$ is represented as a piecewise linear curve consisting of $n^{s}$ uniformly distributed samples $\mathbf{s}_{i,j}$ (with the same distance between two consecutive samples on each strand), along with a corresponding head mesh $\mathcal{M}^{\text{head}}$. The output of our pipeline is a hair card model $\mathcal{M}^{\text{output}}=\langle\{\mathcal{C}_{i}\}_{i=1}^{n^{c}},\{\mathcal{T}_{i}\}_{i=1}^{n^{t}}\rangle$, composed of $n^{c}$ hair cards and $n^{t}$ textures to balance visual fidelity and performance constraints. In order to save texture space, users oftentimes choose $n^{c}>\!\!>n^{t}$ such that multiple hair cards need to share the same texture. A hair card $\mathcal{C}_{i}$ refers to a quad strip with $n^{q}$ consecutive quads. Our objective is to ensure that the synthesized hair card model $\mathcal{M}^{\text{output}}$ approximates the visual appearance of the input $\mathcal{H}$ as closely as possible.
@@ 4.2. Card Geometry Initialization
To initialize the card geometry, we need to build an orthogonal frame system $\{\mathbf{t}_{k,j},\mathbf{n}_{k,j},\mathbf{b}_{k,j}\}$ at each sample $\bar{\mathbf{s}}_{k,j}$ along the central strand $\bar{\mathcal{S}}_{k}$, where $\mathbf{t}_{k,j},\mathbf{n}_{k,j},\mathbf{b}_{k,j}$ are the tangent, normal, and binormal vectors, respectively. The problem of building a frame system for a 3D curve has been thoroughly studied, for which the standard construction is the Frenet-Serret formulas. But this formula suffers from singular configurations and a more numerically stable choice is the Bishop formulas (Bergou et al., 2008), which have been adopted in strand-based hair simulation. Specifically, given the normal $\mathbf{n}_{k,1}$ for the first curve segment, the entire frame system can be computed by solving the Bishop formulas and we refer readers to (Bergou et al., 2008) for the piecewise linear discretization. Given the frame system, we calculate the maximum distance from $\bar{\mathbf{s}}_{k,j}$ to all corresponding strand samples within the cluster along the binormal direction such that:
$W_{k,j}=\max_{\mathcal{S}_{i}\in\mathcal{G}_{k}}\max_{\mathbf{s}_{i,j}\in\mathcal{S}_{i}}\left|(\mathbf{s}_{i,j}-\bar{\mathbf{s}}_{k,j})\cdot\mathbf{b}_{k,j}\right|.$
@@ 4.3. Explicit Hair Card Texture
where $\mathbf{p}_{i,j}^{\bullet}$ are the vertices of the corresponding triangle face in the card geometry that contains $\mathbf{s}_{i,j}^{\Gamma}$, $\mathbf{n}_{i,j}^{\star}$ is the normal of that triangle, and $z\in\mathbb{R}$ is the displacement along $\mathbf{n}_{i,j}^{\star}$. Hence, each hair card texture can be explicitly represented as a set of points $\{\mathbf{s}_{i,j}^{\text{uv}}\}$ embedded within the uv-space. Using $\{\mathbf{s}_{i,j}^{\text{uv}}\}$ as our intermediary representation induces two remarkable features. First, this representation is amenable to end-to-end differentiable rendering, while it does not suffer from aliasing error induced by RGB images. Indeed, for differentiable rendering, 3D hair strands can be first recovered from Eq. 2. These 3D strands are then extended as camera-facing triangle strips with a pre-strand hair width, denoted as $w_{i}$, which is then forwarded to differentiable renderer. Each and every step of this procedure is differentiable and a similar technique is proposed in (Sklyarova et al., 2023). Further, our representation can be converted back to RGB images via rasterization. Finally, to avoid uv-space tangent distortion by the card geometry in world space, we decompose tangent from strand geometry and store it as a per-vertex 3D attribute denoted as $\{\mathbf{t}_{i,j}^{\text{3D}}\}$, which plays a central role in the hair appearance model.
@@ 4.4. Texture Reduction
To facilitate texture sharing and improve both memory efficiency and runtime performance, user could optionally reduce the number of textures before packing them into a texture atlas. Clusters with similar strand counts and shapes are grouped to share the same texture. Specifically, after the per-cluster optimization, we allow users to provide the target texture number $n^{t}$. We then perform another round of k-mean clustering to merge cards with similar textures. To this end, we first rasterize the uv-space strands into RGB textures and then compute the Learned Perceptual Image Patch Similarity (LPIPS) metric (Zhang et al., 2018) between all pairs of card textures. The LPIPS metric is used to guide our k-mean clustering. For each cluster, only the texture closest to the k-mean center is retained, and it is reused across all hair cards within the same cluster. After the clustering, our joint optimization would then fine-tune the shared textures to collectively match the appearance.

### v2

# Auto Hair Card Extraction for Smooth Hair with Differentiable Rendering

@@ Abstract.
Hair cards remain a widely used representation for hair modeling in real-time applications, offering a practical trade-off between visual fidelity, memory usage, and performance. However, generating high-quality hair card models remains a challenging and labor-intensive task. This work presents an automated pipeline for converting strand-based hair models into hair card models with a limited number of cards and textures while preserving the hairstyle appearance. Our key idea is a novel differentiable representation where each strand is encoded as a projected 2D curve in the texture space, which enables end-to-end optimization with differentiable rendering while respecting the structures of the hair geometry. Based on this representation, we develop a novel algorithm pipeline, where we first cluster hair strands into initial hair cards and project the strands into the texture space. We then conduct a two-stage optimization, where our first stage optimizes the orientation of each hair card separately, and after strand projection, our second stage conducts joint optimization over the entire hair card model for fine-tuning. Our method is evaluated on a range of hairstyles, including straight, wavy, curly, and coily hair. To capture the appearance of short or coily hair, our method comes with support for hair caps and cross-card.
@@
copyright: acmlicensedjournal: TOGjournalyear: 2025journalvolume: 44journalnumber: 6article: 1publicationmonth: 12doi: 10.1145/3763295journal: TOGsubmissionid: 1299ccs: Computing methodologies Shape modeling
@@ 1. Introduction
Hair plays a significant role in creating believable characters and immersive environments in films, video games, and virtual reality. However, the average human scalp has approximately 100,000 to 150,000 hair follicles. Although strand-based hair simulation and rendering have become increasingly popular in recent production pipelines due to their ability to realistically capture dynamic behavior and visual richness (Tafuri, 2019; Epic Games, 2021; Huang et al., 2023; Hsu et al., 2023, 2024, 2025), hair cards remain a widely adopted technique in real-time applications (Jiang, 2016). Hair cards are flat, textured quad strips that approximate the visual appearance of hair clusters. As illustrated in Fig. 1, hair cards can effectively reproduce the intricate structure of complex hair styles. Beyond their role in representing high-fidelity hair with a large number of textured planes (Jiang, 2016), hair card representations are also widely used in level-of-detail (LoD), where strand-based models serve as a high-resolution representation, while hair cards can be used as a low-resolution approximation to significantly reduce rendering costs when hair is viewed from a distance. Additionally, mobile games and other low-performance platforms still rely heavily on card-based representations due to strict hardware constraints. Even on high-performance PC platforms, strand-based hair is typically reserved for main characters, while NPCs and background characters use hair cards to reduce runtime costs.
@@
Unfortunately, the industry currently heavily relies on crafting hairstyles manually using hair cards. In this workflow, artists first create a set of hair card textures that contain varying numbers of strands and degrees of curvature. They then manually extrude each hair card from the scalp outward, following the hair flow while varying its length, width, and orientation to achieve a realistic appearance. Apparently, the fidelity quality of the hair card model is primarily bound by the number of cards and textures allowed. As a result, manually crafting a low-resolution hair card model, with a limited number of quads and card textures while maintaining high visual fidelity to the reference strand-based model, is a challenging and labor-intensive task, typically requiring several days to weeks of work from an experienced artist.
@@
Most recent research works have focused on reconstructing strand-based hair models from single-view images (Zhou et al., 2018; Zhang and Zheng, 2019; Wu et al., 2022; Zheng et al., 2023), multi-view inputs (Kuang et al., 2022; Wu et al., 2024b; Zhou et al., 2024), and CT scans (Shen et al., 2023). However, there is a lack of work on converting these reconstructed strands into hair card representations suitable for real-time applications. Extracting hair cards from strand-based hair models presents several challenges. First, there is a strict budget on both the number of hair cards and on the size of texture, imposed by memory and computational efficiency constraints. Therefore, the extraction process must carefully balance visual fidelity with performance. Second, the hairstyle is highly irregular and chaotic, exhibiting significant variation in strand length, shape, and structure, making it challenging to generate a simplified representation while preserving visual details. Recently, Unreal Engine (UE) (Epic Games, 2021) introduced an automatic hair card generation tool from strand-based models (Epic Games, 2025). However, the quality of its generated results remains insufficient for production use, often failing to preserve the appearance of the target hair strand model when the given hair card target is limited.
@@
This paper presents the first automated pipeline that converts straight and wavy strand-based hair models into hair cards. The process begins with hair strand clustering, where strands are grouped based on a hair shape similarity metric. For each cluster, we optimize the orientation of the card geometry to minimize the strand projection error. To improve memory efficiency and runtime performance, we perform second-stage clustering on the hair textures, enabling texture sharing across multiple cards. Finally, we jointly optimize the hair card positions, strand shapes, and hair widths to minimize differences in tangent, depth, and coverage between our output cards and the original strand-based model. To this end, conventional RGB-image texture representations can lead to significant aliasing artifacts. Instead, we propose an explicit hair card representation in which each strand is projected into texture space as a 2D curve. These 2D curves serve as an intermediate representation that enables high-quality rendering and supports differentiable optimization. To enhance the visual fidelity for short and coily hairstyles, which is challenging for original hair card representation, our framework also supports baking a hair cap texture and generating a pair of crossed cards per hair cluster, enriching the volumetric appearance of the hair.
@@
We evaluate our pipeline on a diverse range of hairstyles, including straight, wavy, curly, and coily hair, with a limited number of cards and textures. Experiments show that our automated pipeline surpasses both UE automatic solution and manual-crafted cards.
@@ 2. RELATED WORK
Both hair simulation and rendering are computationally expensive due to the large number of individual strands and their complex interactions. To reduce computational costs, researchers have developed various reduced representations over the years, including 2D strips (commonly referred to as hair cards(Koh and Huang, 2001; Ward et al., 2003), cubic lattice structures(Volino and Magnenat-Thalmann, 2006), short hair strips (Guang and Zhiyong, 2002), and volumetric representations (Wu and Yuksel, 2016; Lee et al., 2019). During rendering, these reduced representations are expanded into full hair using baked textures or procedural functions. Hair cards remain widely used in the gaming industry due to their simplicity and efficiency (Jiang, 2016). For a comprehensive overview of hair rendering and simulation, we refer readers to the course by Bertails et al. (2008). However, creating hair cards for multiple LoDs continues to be a labor-intensive process for artists, representing a significant bottleneck in production pipelines. Recently, Huang et al. (2025) presents a real-time framework to create and render hair LoD dynamically based on pre-clustered hairs.
@@
Traditionally, artists have used tools like Maya XGen to create strand-based and card-based hair models. However, manually crafting hair remains a labor-intensive process. To simplify hair authoring, Yuksel et al. (2009) introduced hair mesh, a volumetric representation that provides high-level editing tools for artists. Sketch-based input methods (Fu et al., 2007; Shen et al., 2020) have also offered user-friendly interaction capability, allowing artists to design hair directly on the screen. For automatic hair modeling, researchers have explored extracting strand-based geometries directly from images. Earlier methods relied on heuristic-based approaches (Kong and Nakajima, 1998; Paris et al., 2008; Jakob et al., 2009; Sun et al., 2021; Hu et al., 2017) or leveraged large 3D hair databases (Hu et al., 2015; Chai et al., 2016; Liang et al., 2018) for guidance. Recently, learning-based techniques have demonstrated accuracy and robustness in reconstructing simple hairstyles from single-view images (Zhou et al., 2018; Zhang and Zheng, 2019; Wu et al., 2022; Zheng et al., 2023) and multi-view inputs (Kuang et al., 2022; Rosu et al., 2022; Sklyarova et al., 2023; Wu et al., 2024b; Takimoto et al., 2024; Zakharov et al., 2024; Zhou et al., 2024). Wu et al. (2024a) present a geometric method to generate highly coiled hair. Beyond explicit modeling and reconstruction, hair synthesis methods enable hairstyle transfer from one to another using feature maps (Wang et al., 2009), further refined with learning-based features (Zhou et al., 2023; Sklyarova et al., 2024; Chen et al., 2024; He et al., 2025). However, these techniques primarily aim to generate explicit strand geometries, which are computationally intensive and unsuitable for real-time applications. There are also learning-based methods using neural representations (Luo et al., 2024; Zheng et al., 2025; Wang et al., 2023). While neural-based methods have shown significant progress in reconstructing and representing complex hair geometry, integrating these representations into industrial game engines in a highly efficient manner remains a major challenge.
@@
In addition to hair, arbitrarily oriented billboards (or impostors) with textures are widely used for rendering trees and forests (Behrendt et al., 2005) and large-scale scenes (Lall et al., 2018; Hladky et al., 2022). To convert a 3D mesh to billboards, Décoret et al. (2003) proposed a greedy optimization algorithm that approximates the 3D model using a discretized plane parameterization in spherical coordinates. This method was later enhanced by Andujar et al. (2004) to produce better-fitting billboards. Silvennoinen et al. (2014) uses a similar concept to generate sets of planars as occluders. However, as hair presents unique challenges due to its extensive strand number and complex geometry, existing billboard generation methods cannot be applied directly.
@@ 3. Problem Statement and Overview
The input to our method is a strand-based hairstyle model $\mathcal{H}=\{\mathcal{S}_{i}\}$, where each strand $\mathcal{S}_{i}=\{\mathbf{s}_{i,j}\}$ is represented as a piecewise linear curve consisting of $n^{s}$ uniformly distributed samples $\mathbf{s}_{i,j}$ (with the same distance between two consecutive samples on each strand), along with a corresponding head mesh $\mathcal{M}^{\text{head}}$. The output of our pipeline is a hair card model $\mathcal{M}^{\text{output}}=\langle\{\mathcal{C}_{i}\}_{i=1}^{n^{c}},\{\mathcal{T}_{i}\}_{i=1}^{n^{t}}\rangle$, composed of $n^{c}$ hair cards and $n^{t}$ textures to balance visual fidelity and performance constraints. In order to save texture space, users oftentimes choose $n^{c}>\!\!>n^{t}$ such that multiple hair cards need to share the same texture. A hair card $\mathcal{C}_{i}$ refers to a quad strip with $n^{q}$ consecutive quads. Our objective is to ensure that the synthesized hair card model $\mathcal{M}^{\text{output}}$ approximates the visual appearance of the input $\mathcal{H}$ as closely as possible.
@@ 4.2. Card Geometry Initialization
To initialize the card geometry, we need to build an orthogonal frame system $\{\mathbf{t}_{k,j},\mathbf{n}_{k,j},\mathbf{b}_{k,j}\}$ at each sample $\bar{\mathbf{s}}_{k,j}$ along the central strand $\bar{\mathcal{S}}_{k}$, where $\mathbf{t}_{k,j},\mathbf{n}_{k,j},\mathbf{b}_{k,j}$ are the tangent, normal, and binormal vectors, respectively. The problem of building a frame system for a 3D curve has been thoroughly studied, for which the standard construction is the Frenet-Serret formulas. But this formula suffers from singular configurations, and a more numerically stable choice is the Bishop formulas (Bergou et al., 2008), which have been adopted in strand-based hair simulation. Specifically, given the normal $\mathbf{n}_{k,1}$ for the first curve segment, the entire frame system can be computed by solving the Bishop formulas, and we refer readers to (Bergou et al., 2008) for the piecewise linear discretization. Given the frame system, we calculate the maximum distance from $\bar{\mathbf{s}}_{k,j}$ to all corresponding strand samples within the cluster along the binormal direction such that:
(1) $W_{k,j}=\max_{\mathcal{S}_{i}\in\mathcal{G}_{k}}\left|(\mathbf{s}_{i,j}-\bar{\mathbf{s}}_{k,j})\cdot\mathbf{b}_{k,j}\right|.$
@@ 4.3. Explicit Hair Card Texture
where $\mathbf{p}_{i,j}^{\bullet}$ are the vertices of the corresponding triangle face in the card geometry that contains $\mathbf{s}_{i,j}^{\Gamma}$, $\mathbf{n}_{i,j}^{\star}$ is the normal of that triangle, and $z\in\mathbb{R}$ is the displacement along $\mathbf{n}_{i,j}^{\star}$. Hence, each hair card texture can be explicitly represented as a set of points $\{\mathbf{s}_{i,j}^{\text{uv}}\}$ embedded within the uv-space. Using $\{\mathbf{s}_{i,j}^{\text{uv}}\}$ as our intermediary representation induces two remarkable features. First, this representation is amenable to end-to-end differentiable rendering, while it does not suffer from aliasing error induced by RGB images. Indeed, for differentiable rendering, given the card mesh and explicit hair card texture, 3D hair strands can be first recovered from Eq. 3. The user-defined per-strand hair width is denoted as $w_{i}$. Each and every step of this procedure is differentiable, and a similar technique is proposed in (Sklyarova et al., 2023). Further, our representation can be converted back to RGB images via rasterization. Finally, to avoid uv-space tangent distortion by the card geometry in world space, we decompose the tangent from strand geometry and store it as a per-vertex 3D attribute denoted as $\{\mathbf{t}_{i,j}^{\text{3D}}\}$, which plays a central role in the hair appearance model. The renderer can also produce depth and coverage maps by utilizing depth and flat color as attributes for each vertex.
@@ 4.4. Texture Reduction
To facilitate texture sharing and improve both memory efficiency and runtime performance, users could optionally reduce the number of textures before packing them into a texture atlas. Clusters with similar strand counts and shapes are grouped to share the same texture. Specifically, after the per-cluster optimization, we allow users to provide the target texture number $n^{t}$. We then perform another round of k-means clustering to merge cards with similar textures. To this end, we first rasterize the UV-space strands into RGB textures and then compute the Learned Perceptual Image Patch Similarity (LPIPS) metric (Zhang et al., 2018) between all pairs of card textures. The LPIPS metric is used to guide our k-means clustering. For each cluster, only the texture closest to the k-mean center is retained, and it is reused across all hair cards within the same cluster. After the clustering, our joint optimization would then fine-tune the shared textures to collectively match the appearance.